• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name jpchen

    Log entry time 18:37:59 on November 25, 2008

    Entry number 250720

    keyword=Target position survey

    According to Joe, who talked to the Survey group after the target survey, the target position offsets are still the same as before. Yaw was reduced. We will start our run with the same beam positions on BPM A (1.0,1.3) and BPMB (-0.5, 1.0) as before this down.